### v3

@@ 4.2 Choice of tuning parameters for polynomial-regression FAT
${y}_{it}^{(1)}(0)$ $=\mu_{i}+\rho{y}_{it-1}^{(1)}(0)+u_{it},\;t\geq 1,{y}_{i0}^{(1)}(0)\sim\mathcal{N}\left(\frac{\mu_{i}}{1-\rho},\frac{1}{1-\rho^{2}}\right),$
${y}_{it}^{(2)}(0)$ $={y}_{it-1}^{(2)}(0)+\epsilon_{it},\;t\geq 1,{y}_{i0}^{(2)}(0)=0,$
@@
$\mu_{i}$ $\sim\mathcal{U}\left[-1,1\right],\;u_{it}\sim\mathcal{N}\left(0,1\right),\;\epsilon_{it}\sim\mathcal{N}(0,1),$
$\rho$ $=0.2,\;\delta_{i}=1,T=6,\;\tau=5,\;h=1,\;n=1000.$
@@
Table 2 shows results for the bias and standard error of $\widehat{\rm FAT}^{\rm PR}_{1}$ across different tuning parameters. The table shows that, when the DGP is mean stationary (first panel) or when it is the sum of a mean stationary and a random walk (second panel), the bias does not vary much across different values of the tuning parameters, while the standard errors are smaller for smaller $q$ and larger $R$. When the DGP contains a time trend component, we observe bias when the polynomial-order $q$ is less than the true order of the time trend, as the theory predicts. In this case, however, a smaller estimation window $R$ gives a smaller bias. When $q\geq 1$, the performance of the estimator in terms of bias is again robust to the choice of tuning parameters, with smaller standard errors for smaller $q$ and larger $R$.
@@
Table 2: Bias and standard error (s.e.) for the polynomial-regression FAT when the counterfactual is specified as indicated in the left-most column.
@@ 5 Empirical illustration
We illustrate how our approach without a control group could reach similar conclusions as existing approaches that are based on the availability of a control group. [^30] Additional replications are included in the Online Appendix C.4 and C.5. We consider the application in Goodman-Bacon (2021) on the effect of no-default laws on female suicide. The data is on U.S. states that adopted no-default divorce laws from 1969 to 1985. The outcome is the age-adjusted average suicide mortality rate per million women (ASMR) in the state. We work with a balanced panel of $n=36$ treated states and 10 time periods, 5 of which are pre-treatment. [^31] All treated states have at least 5 periods of pre-treatment data, which is why we consider this particular balanced panel.
We assume the same tuning parameters $q$ and $R$ for FAT across states. To help select the basis functions and the tuning parameters, Figure 1 plots the outcome variable averaged across states, as a function of time to adoption. The visible trending behaviour (plausibly reflecting a deterministic component) in the pre-treatment data suggests using a polynomial basis function and choosing orders $q>0$. There is no strong indication of structural instability in the pre-treatment data, so we use all available data for estimation, that is, we set $R=5$. We compute the polynomial-regression based forecast, $\widehat{y}^{PR}_{i\tau+h}$ in (29), for a number of forecast horizons $h$ and orders $q$. FAT estimates at $\tau+h$ are shown in the top panel of Figure 2 as a function of $h\in\{1,\dots,5\}$ and for $q\in\{1,2\}$. The variance of the estimator is computed as $\frac{1}{n}\sum_{i=1}^{n}(y_{i\tau+h}-\widehat{y}^{PR}_{i\tau+h}-\widehat{{\rm FAT}}_{h}^{PR})^{2}$. The results show a statistically insignificant decrease in the suicide rate after adoption of the no-fault divorce laws across different forecast horizons. This replicates the findings in Goodman-Bacon (2021), Clarke and Tapia-Schythe (2021), but without using a control group. As expected, compared to the results in Figure 2 in Clarke and Tapia-Schythe (2021), our $95\%$ confidence intervals are greater for higher forecast horizons, e.g., $h\geq 4$.
To validate the procedure, we compute placebo FATs. In the bottom panel of Figure 2, we report placebo FAT at lag $j$, $j\in\{0,\dots,3\}$, calculated as if the law was adopted $j$ years earlier than its actual adoption date (for each state). The forecast horizon is $h=1$, so that FAT at lag 0 here corresponds to FAT at horizon 1 in the upper panel. The figure shows that the placebo FATs for both polynomial orders are statistically insignificant. This offers suggestive evidence that FAT estimates may be interpreted as ATT estimates.
@@
Figure 2: The top panel shows estimates of FAT for different forecast horizons (horizontal axis) and polynomial orders (different colors). The bottom panel shows estimates of placebo FAT (for $h=1$) as a function of the lag (horizontal axis) and polynomial order (different lines). Lag 0 is the actual adoption-year and placebo FAT at lag 0 is the FAT at forecast horizon 1 in the top panel. The error bars in both panels represent $95\%$ confidence intervals.
@@ 6 Conclusion
This paper proposed an estimator of the average treatment effects in the absence of a control group, based on forecasting individual counterfactuals using basis function regressions over a (short) time series of pre-treatment data. Forecast unbiasedness is a key requirement that is satisfied by our approach under a broad class of DGPs that express the individuals counterfactuals as the sum of up to three unobserved components: a stationary process, a stochastic trend, and a deterministic trend. The approach is robust, general and flexible, allowing for unbalanced panels, heterogeneous treatment effects, and staggered treatment timing. Forecasting counterfactuals using a model instead requires stronger assumptions and can perform poorly due to misspecification and estimation bias in small samples (e.g., incidental parameter problem).
@@ Appendix A Proofs
$X_{i}\equiv\left[\begin{array}[]{cccc}1&b_{1}\left(c_{1}\right)&\dots&b_{q_{i}}\left(c_{1}\right)\\ 1&b_{1}\left(c_{2}\right)&\dots&b_{q_{i}}\left(c_{2}\right)\\ 1&b_{1}\left(c_{3}\right)&\dots&b_{q_{i}}\left(c_{3}\right)\\ \dots&\dots&\dots&\dots\\ 1&b_{1}\left(c_{\tau_{i}}\right)&\dots&b_{q_{i}}\left(c_{\tau_{i}}\right)\end{array}\right],\text{}H_{i}\equiv\left[\begin{array}[]{cccc}1&b_{1}\left(\tau+h\right)&\dots&b_{q_{i}+1}\left(\tau+h\right)\end{array}\right].$ (31)
@@
$\hat{\alpha}^{\left(q_{i},R_{i}\right)}=\left(X_{i}^{\prime}X_{i}\right)^{-1}X_{i}^{\prime}y_{i},$
@@ Appendix B Estimating the variance of $\widehat{{\rm FAT}}_{h}$
If an additional unknown common parameter $\beta_{0}$ needs to be estimated, then the uncertainty of the estimator $\widehat{\beta}$ also becomes relevant. In that case, according to Theorem 3, the asymptotic variance of $\sqrt{n}\widehat{{\rm FAT}}_{h}$ is $\bar{\sigma}^{*2}_{n}:={\rm Var}\left(\frac{1}{\sqrt{n}}\sum_{i}u^{*}_{i\tau+h}\right)$, where $u^{*}_{i\tau+h}:=y_{i\tau+h}-\widehat{y}_{h}(\beta_{0},y_{i},x_{i})-\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}\left[\frac{\partial\widehat{y}_{h}(\beta_{0},y_{j},x_{j})}{\partial\beta^{\prime}}\right]\psi(y_{i},x_{i})$. Here, both $\beta_{0}$ and the influence function $\psi(y_{i},x_{i})$ of $\widehat{\beta}$ are unknown. For example, if $\widehat{\beta}$ is a method of moment estimator (which includes OLS, MLE, and exactly identified IV) that solves the sample moment condition
@@
$\widehat{u}^{*}_{i\tau+h}=y_{i\tau+h}-\widehat{y}_{h}(\widehat{\beta},y_{i},x_{i})-\left(\frac{1}{n}\sum_{j=1}^{n}\ \frac{\partial\widehat{y}_{h}(\beta_{0},y_{j},x_{j})}{\partial\beta^{\prime}}\right)\widehat{\psi}(\widehat{\beta},y_{i},x_{i}),$
@@ Appendix C Online Appendix
The Online Appendix contains extensions, additional simulations and empirical replications. In the main text, we consider FAT under the assumption that the class of DGPs for the counterfactuals does not include common unforecastable shocks. We show how this assumption can be relaxed when there exists another variable that is subject to the same common shock but not to the treatment (in Online Appendix C.1) or when there exists a control group that is subject to the common shock (in Online Appendix C.2). The simulations in the main text consider the case of homogeneous time trends and autoregressive parameters. We relax this in Online Appendix C.3. Finally, we present two additional empirical applications in Online Appendix C.4 and C.5.
C.1 Extension: no control group, common shock
We sketch here a possible way to weaken the assumption of no forecastable shock between treatment time $\tau$ and forecast time $\tau+h,h\geq 1$.
Let $\widetilde{y}_{it}\left(0\right)$ follow either Assumption 3 or 4 and let $W_{it}$ be another observed variable that is affected by a shock $\gamma_{t}$ but not by the treatment. Suppose that
$W_{it}$ $=\theta_{i}^{W}\gamma_{t}1\left(t>\tau\right)+\epsilon_{it}.$ (35)
Denote by $\widehat{\gamma}_{t}$ the estimator of $\gamma_{t}$ obtained from post-treatment data on $W_{it},$ $t>\tau$ via, e.g., PCA. Then
where $\widehat{\widetilde{y}}_{i\tau+h}(0)$ is the couterfactual forecast obtained as before via basis function regression.
C.2 Extension: control group
@@
Without a control group, Section 2 derived sufficient conditions ensuring that ${\rm FAT}_{h}$ defined in (6) is a consistent and asymptotically normal estimator of ${\rm ATT}_{h}$ defined in (4). These conditions are the ability to obtain forecasts of the counterfactuals using pre-treatment data that are on average unbiased (Assumption 1) and the validity of a central limit theorem (Assumption 2). As discussed in the main text, these conditions exclude the presence of time effects such as macro shocks that affect all individuals between times $\tau$ and $\tau+h$, $h\geq 1$, and that are unforecastable using pre-treatment data. The presence of a control group allows us to weaken this assumption.
C.2.1 DFAT: FAT with a control group
Suppose that all individuals are untreated before the implementation of the treatment at time $\tau$ and that some individuals remain untreated after $\tau$. Let $D_{i}=1$ if individual $i$ is untreated before and after $\tau$. The observed outcome of individual $i$ at time $t$ is then
@@ C.2.1 DFAT: FAT with a control group
where $n_{1}$ is the number of treated individuals at time $\tau+h$, $n_{0}$ is the number of control individuals at time $\tau+h$, and $y_{i\tau+h}$ is the observed outcome at $\tau+h$ given by (37).
@@
Unlike in the baseline case, the forecast $\widehat{y}_{i\tau+h}(0)$ can be biased, as long as the average bias for the treated group equals the average bias for the control group. As a consequence, the DGP for $y_{it}(0)$ can contain additive time effects that are common across individuals such as additive macro shocks that affect both treated and control groups in the same way.
The presence of a control group allows us to substitute Assumption 4 to allow for a common shock that is not necessarily polynomial, e.g., $y_{it}\left(0\right)=\tilde{y}_{it}(0)+\gamma_{t}$, where $\tilde{y}_{it}(0)$ satisfies Assumption 4. Then, under assumptions similar to Assumptions 1 and 2, it is possible to show consistency and asymptotic normality of (39).
As in Section 2 in the main text, we suggest using $\widehat{y}_{i\tau+h}^{(q_{i},R_{i})}(0)$ as an estimator for $\widehat{y}_{i\tau+h}(0)$ in (39). Here, the parameters $q_{i},R_{i}$ do not necessarily have to be the same for the treated and control units.
C.2.2 Comparison with Difference-in-Differences
Despite the apparent similarity with the difference-in-differences (DiD) framework, our method in the presence of a control group allows for DGPs for counterfactuals with more general forms of latent heterogeneity. For example, our approach allows for counterfactuals that follow fully heterogeneous autoregressive processes and/or unit root processes. In addition, it allows for the DGPs to have additive individual-specific time trends, as long as the deterministic time trend is either known or can be approximated by, e.g., a polynomial.
@@ C.2.2 Comparison with Difference-in-Differences
$y_{it}\left(0\right)=\rho y_{it-1}\left(0\right)+\gamma_{t}+k_{i}t+\epsilon_{it},\mathbb{E}\left(\epsilon_{it}\right)=0,$
where $\rho\in[0,1]$, $\gamma_{t}$ is a common shock and $k_{i}$ is an individual-specific time trend coefficient.
@@
$\frac{1}{n}\sum^{n}_{i=1}\mathbb{E}(\left.y_{i0}(0)\right|D_{i}=1)=\frac{1}{n}\sum^{n}_{i=1}E(\left.y_{i0}(0)\right|D_{i}=0)$
@@
[…]
C.2.3 Related literature
[…]
C.3 Simulation: Heterogeneous coefficients
We compare the finite-sample behavior of the polynomial-regression FAT when the counterfactual process satisfies Assumption 4 with heterogeneous coefficients $\rho_{i}$ and $\delta_{i}$. That is, we consider the same specification of $y_{it}(0)$ as in Section 4.2 with $I_{1}=I_{2}=I_{3}=1$, with the only changes being that $\rho_{i}$ and $\delta_{i}$ vary across individuals.

### v4

@@ 4.2 Choice of tuning parameters for Unobserved-Components FAT
${y}_{it}^{(1)}(0)$ $=\mu_{i}+\rho{y}_{it-1}^{(1)}(0)+u_{it},\;t\geq 1,\qquad{y}_{i0}^{(1)}(0)\sim\mathcal{N}\left(\frac{\mu_{i}}{1-\rho},\frac{1}{1-\rho^{2}}\right),$
${y}_{it}^{(2)}(0)$ $={y}_{it-1}^{(2)}(0)+\epsilon_{it},\;t\geq 1,\qquad{y}_{i0}^{(2)}(0)=0,$
@@
$\mu_{i}$ $\sim\mathcal{U}\left[-1,1\right],\quad u_{it}\sim\mathcal{N}\left(0,1\right),\quad\epsilon_{it}\sim\mathcal{N}(0,1),$
$\rho$ $=0.2,\quad\delta_{i}=1,\quad T=6,\quad\tau=5,\quad h=1,\quad n=1000.$
@@
Table 5 shows results for the bias and standard error of $\widehat{\rm FAT}^{\rm UC}_{1}$ across different tuning parameters. The table shows that, when the DGP is mean stationary (first panel) or when it is the sum of a mean stationary and a random walk (second panel), the bias does not vary much across different values of the tuning parameters, while the standard errors are smaller for smaller $q$ and larger $R$. When the DGP contains a time trend component, we observe bias when the polynomial-order $q$ is less than the true order of the time trend, as the theory predicts. In this case, however, a smaller estimation window $R$ gives a smaller bias. When $q\geq 1$, the performance of the estimator in terms of bias is again robust to the choice of tuning parameters, with smaller standard errors for smaller $q$ and larger RR. [^15] We observe the same behavior with a sample size of $N=30$, see Table 8 in the Online Appendix.
@@
Table 5: Bias and standard error (s.e.) for the Unobserved-Components FAT when the counterfactual is specified as indicated in the left-most column. Sample size $N=1000$.
@@ 5 Empirical illustration
We consider an empirical exercise showing that FAT, despite not requiring an untreated control group, yields conclusions consistent with established methods that rely on a control group. This positions FAT as a valuable complementary robustness check when a control group is available, rather than a replacement for existing methods such as TWFE or DiD. [^16] Additional replications are included in the Online Appendix C.3 and C.4.
We consider the application in Goodman-Bacon (2021) on the effect of no-fault laws on female suicide. The data are on U.S. states that adopted no-fault divorce laws from 1969 to 1985. The outcome is the age-adjusted average suicide mortality rate per million women (ASMR) in the state. We work with a balanced panel of $n=36$ treated states and 10 time periods, 5 of which are pre-treatment.
We assume the same tuning parameters $q$ and $R$ for FAT across states. To help select the basis functions and the tuning parameters, Figure 1 plots the outcome variable averaged across states, as a function of time to adoption. The visible trending behaviour (plausibly reflecting a deterministic component) in the pre-treatment data suggests using a polynomial basis function and choosing orders $q>0$. There is no strong indication of structural instability in the pre-treatment data, so we use all available data for estimation, that is, we set $R=5$. We compute the forecast $\widehat{y}^{UC}_{i\tau+h}$ in (29) for a number of forecast horizons $h$ and orders $q$. The Unobserved-Components FAT estimates at $\tau+h$ are shown in the top panel of Figure 2 as a function of $h\in\{1,\dots,5\}$ and for $q\in\{1,2\}$. The variance of the estimator is computed as explained in Section B. The results show a statistically insignificant decrease in the suicide rate after adoption of the no-fault divorce laws across different forecast horizons. This replicates the findings in Goodman-Bacon (2021), Clarke and Tapia-Schythe (2021), but without using a control group. The standard error of our estimator is comparable to that obtained in the above-mentioned studies that employ event-study design approaches and, as expected, our $95\%$ confidence intervals are greater for longer forecast horizons (see, e.g., the results in Figure 2 in Clarke and Tapia-Schythe 2021). [^17] To see how our estimator performs in terms of standard error with a small sample size, as in this application, see Table 8 in the Online Appendix for simulation results with $N=30$.
To validate the procedure, in Figure 2 we report placebo FAT at lag $j$, $j\in\{0,\dots,3\}$, calculated as if the law was adopted $j$ years earlier than its actual adoption date (for each state). The forecast horizon is $h=1$. The figure shows that the placebo FATs for both polynomial orders are close to zero. This offers suggestive evidence that FAT estimates may be interpreted as ATT estimates.
@@
Figure 2: FAT estimates for different forecast horizons (horizontal axis) and polynomial orders (different colors). To the left of the dashed vertical line, the figure shows placebo FAT (for $h=1$) as a function of the lag (horizontal axis) and polynomial order (different lines). The error bars in both panels represent $95\%$ confidence intervals.
@@ 6 Conclusion
This paper proposes an estimator of the average treatment effects in the absence of an untreated or control group, based on forecasting individual counterfactuals using basis function regressions over a (short) time series of pre-treatment data. Forecast unbiasedness is a key requirement that is satisfied by our approach under a broad class of DGPs that express the individual counterfactuals as the sum of up to three unobserved components: a stationary process, a stochastic trend, and a deterministic trend. The approach is robust, general, and flexible, allowing for unbalanced panels, heterogeneous treatment effects, and staggered treatment timing. Forecasting counterfactuals using a parametric model instead requires stronger assumptions and can perform poorly due to misspecification and estimation bias in small samples (e.g., incidental parameter problem).
@@ Appendix A Proofs
$X_{i}\equiv\left[\begin{array}[]{cccc}1&b_{1}\left(c_{1}\right)&\dots&b_{q_{i}}\left(c_{1}\right)\\ 1&b_{1}\left(c_{2}\right)&\dots&b_{q_{i}}\left(c_{2}\right)\\ 1&b_{1}\left(c_{3}\right)&\dots&b_{q_{i}}\left(c_{3}\right)\\ \dots&\dots&\dots&\dots\\ 1&b_{1}\left(c_{R_{i}}\right)&\dots&b_{q_{i}}\left(c_{R_{i}}\right)\end{array}\right],\text{\;}H_{i}\equiv\left[\begin{array}[]{cccc}1&b_{1}\left(\tau+h\right)&\dots&b_{q_{i}}\left(\tau+h\right)\end{array}\right].$ (32)
@@
$\widehat{c_{i}}^{(q_{i},R_{i})}=\left(X_{i}^{\prime}X_{i}\right)^{-1}X_{i}^{\prime}y_{i},$
@@ Appendix B Estimating the variance of $\widehat{{\rm FAT}}_{h}$
If an additional unknown common parameter $\beta_{0}$ needs to be estimated, the uncertainty of the estimator $\widehat{\beta}$ also becomes relevant. In that case, according to Theorem 3, the asymptotic variance of $\sqrt{n}\widehat{{\rm FAT}}_{h}$ is $\bar{\sigma}^{*2}_{n}:={\rm Var}\left(\frac{1}{\sqrt{n}}\sum_{i}u^{*}_{i\tau+h}\right)$, where $u^{*}_{i\tau+h}:=y_{i\tau+h}-\widehat{y}_{h}(\beta_{0},y_{i},x_{i})-\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}\left[\frac{\partial\widehat{y}_{h}(\beta_{0},y_{j},x_{j})}{\partial\beta^{\prime}}\right]\psi(y_{i},x_{i})$. Here, both $\beta_{0}$ and the influence function $\psi(y_{i},x_{i})$ of $\widehat{\beta}$ are unknown. For example, if $\widehat{\beta}$ is a method of moment estimator (which includes OLS, MLE, and exactly identified IV) that solves the sample moment condition
@@
$\widehat{u}^{*}_{i\tau+h}=y_{i\tau+h}-\widehat{y}_{h}(\widehat{\beta},y_{i},x_{i})-\left(\frac{1}{n}\sum_{j=1}^{n}\ \frac{\partial\widehat{y}_{h}(\widehat{\beta},y_{j},x_{j})}{\partial\beta^{\prime}}\right)\widehat{\psi}(\widehat{\beta},y_{i},x_{i}),$
@@ Appendix C Online Appendix
The Online Appendix contains extensions, additional simulations and empirical replications. We first discuss how to extend the procedure when there exists an untreated group (in Online Appendix C.1). The simulations in the main text consider the case of homogeneous time trends and autoregressive parameters. We relax this in Online Appendix C.2. Finally, we present two additional empirical applications in Online Appendix C.3 and C.4.
C.1 Extension: untreated group
@@
Without an untreated group, Section 2 derived sufficient conditions ensuring that ${\rm FAT}_{h}$ defined in (6) is a consistent and asymptotically normal estimator of ${\rm ATT}_{h}$ defined in (4). These conditions are the ability to obtain forecasts of the counterfactuals using pre-treatment data that are on average unbiased (Assumption 1) and the validity of a central limit theorem (Assumption 2). As discussed in the main text, these conditions exclude the presence of time effects such as macro shocks that affect all individuals between times $\tau$ and $\tau+h$, $h\geq 1$, and that are unforecastable using pre-treatment data. The presence of an untreated group allows us to weaken this assumption.
C.1.1 DFAT: FAT with an untreated group
Suppose that all individuals are untreated before the implementation of the treatment at time $\tau$ and that some individuals remain untreated after $\tau$. Let $D_{i}=1$ if individual $i$ is treated after $\tau$. The observed outcome of individual $i$ at time $t$ is then
@@ C.1.1 DFAT: FAT with an untreated group
where $n_{1}$ is the number of treated individuals at time $\tau+h$, $n_{0}$ is the number of untreated individuals at time $\tau+h$, and $y_{i\tau+h}$ is the observed outcome at $\tau+h$ given by (35).
@@
Unlike in the baseline case, the forecast $\widehat{y}_{i\tau+h}(0)$ can be biased, as long as the average bias for the treated group equals the average bias for the untreated group. As a consequence, the DGP for $y_{it}(0)$ can contain additive time effects that are common across individuals such as additive macro shocks that affect both treated and untreated groups in the same way.
Condition (38) is satisfied when the DGP components in Assumption 4 have the same average conditional expectation across treatment groups, i.e., when the mean-stationary, unit-root, and deterministic trend components are balanced on average between treated and untreated units. This is distinct from parallel trends: parallel trends restricts the evolution of outcome differences over time, while our condition restricts balance in the components that determine forecast errors.
The presence of an untreated group allows us to substitute Assumption 4 to allow for a common shock that is not necessarily polynomial, e.g., $y_{it}\left(0\right)=\tilde{y}_{it}(0)+\gamma_{t}$, where $\tilde{y}_{it}(0)$ satisfies Assumption 4. Then, under assumptions similar to Assumptions 1 and 2, it is possible to show consistency and asymptotic normality of (37).
As in Section 2 in the main text, we suggest using $\widehat{y}_{i\tau+h}^{(q_{i},R_{i})}(0)$ as an estimator for $\widehat{y}_{i\tau+h}(0)$ in (37). Here, the parameters $q_{i},R_{i}$ do not necessarily have to be the same for the treated and untreated units.
C.1.2 Comparison with Difference-in-Differences
Despite the apparent similarity with the difference-in-differences (DiD) estimator, our method in the presence of an untreated group allows for DGPs for counterfactuals with more general forms of latent heterogeneity and outcome dynamics. For example, our approach allows for counterfactuals that follow fully heterogeneous autoregressive processes and/or unit root processes. In addition, it allows for the DGPs to have additive individual-specific time trends, as long as the deterministic time trend is either known or can be approximated by, e.g., a polynomial.
@@ C.1.2 Comparison with Difference-in-Differences
$y_{it}\left(0\right)=\rho y_{it-1}\left(0\right)+\gamma_{t}+k_{i}t+\epsilon_{it},\quad\mathbb{E}\left(\epsilon_{it}\right)=0,$
where $\rho\in[0,1]$, $\gamma_{t}$ is a common shock, and $k_{i}$ is an individual-specific time trend coefficient.
@@
$\frac{1}{n}\sum^{n}_{i=1}\mathbb{E}(\left.y_{i0}(0)\right|D_{i}=1)=\frac{1}{n}\sum^{n}_{i=1}\mathbb{E}(\left.y_{i0}(0)\right|D_{i}=0),$
@@
Remark (Dynamics and Transformed Outcomes). It is theoretically possible to reconcile outcome dynamics with parallel trends by applying the parallel trends assumption to a transformed outcome. As emphasized by recent literature (and discussed in Marx et al. (2024) in the context of sequential exchangeability), if the untreated outcome follows a dynamic process
$Y_{it}(0)=\alpha_{i}+\rho Y_{i,t-1}(0)+\lambda_{t}+u_{it},\quad\mathbb{E}[u_{it}]=0,$
one can define a transformed variable $w_{it}(\rho)\equiv Y_{it}-\rho Y_{i,t-1}$. If parallel trends hold for $w_{it}(\rho)$—i.e., $\mathbb{E}[w_{it}(\rho)-w_{is}(\rho)\mid D_{i}]=\mathbb{E}[w_{it}(\rho)-w_{is}(\rho)]$—then a standard DiD estimator applied to this transformed outcome identifies the ATT. Given enough pre-treatment periods, $\rho$ can be estimated consistently via Anderson-Hsiao.
However, the class of data generating processes handled by the FAT framework is broader in an important sense. The dynamic DiD specification above, as well as standard GMM dynamic-panel estimators (Arellano and Bond, 1991; Blundell and Bond, 1998), typically require a homogeneous autoregressive coefficient $\rho$. Even in flexible short-panel frameworks like Arellano and Bonhomme (2012), the coefficient on the lagged outcome is often restricted. In contrast, FAT accommodates models of the form
$Y_{it}(0)=\rho_{i}Y_{i,t-1}(0)+\text{trend}_{i}+U_{it},$
where the autoregressive persistence $\rho_{i}$ varies across units. These types of models fall outside the scope of most existing short-panel estimators, which typically require restrictions on the form of dynamic heterogeneity.
In addition, our paradigm of first obtaining individual-specific forecasts and only afterwards averaging across individuals avoids any concern about estimating weighted as opposed to unweighted treatment effects. It is well known that for unbalanced panels and for staggered adoption designs the DiD method will estimate a weighted average of the individual specific treatment effects with weights that are determined implicitly by the regression design, and that may even be negative in some cases (see e.g. de Chaisemartin and D’Haultfoeuille 2020; Goodman-Bacon 2021). By contrast, FAT (or DFAT) explicitly gives a weight of $1/n$ (or $1/n_{d}$) to each individual by construction, in accordance with the unweighted treatment effect that is specified as estimand.
C.1.3 Related literature
Conceptually, when there exists an untreated group, our solution resembles difference-in-differences or, more generally, an “event-study design analysis”, e.g., Borusyak et al. (2023). Although the estimated outcome equations may look similar, there is an important distinction between these methods and ours. For example, the extension of FAT to the case of a control group (which we call DFAT) uses control groups to correct for the effect of a common shock, whereas the other methods use control groups to correct for selection into treatment (under different assumptions). Additionally, FAT allows for heterogeneous nonlinear time trends as well as for heterogeneous effects of lagged pre-treatment outcomes. In contrast, there is no straightforward way to control for pre-treatment lagged outcomes in the specifications of, e.g., Sun and Abraham (2020); Callaway and Sant’Anna (2021); Borusyak et al. (2023). In fact, recent literature shows that in the presence of outcome dynamics, conventional estimators used in this literature produce biased estimates for the treatment effect (Botosaru and Liu, 2025; Cornwall and Wentland, 2025).
The synthetic control method has been increasingly used to evaluate the effect of interventions implemented at an aggregate level, see Abadie (2021) for a recent review. In the conventional setting for synthetic controls, there is only one unit that is treated and there are many untreated units that could be used as pseudo-controls, i.e. these are untreated units selected such that the weighted average of their past outcomes “resembles” the trajectory of past outcomes of the treated unit. The counterfactual outcome for the treated unit is then constructed as a weighted average of the post-treatment outcomes of the selected pseudo-control units. In comparison, in our baseline setting, all individuals in the population are treated and there are no control units. The counterfactual outcome for each treated unit is a weighted average of the unit’s own past outcomes. The properties of our estimator rest on averaging across many treated units, an advantage of which is standard inference. Our results apply even when the number of pre-treatment time periods is small, and we fully characterize the class of DGPs that obtains a consistent estimator of the ATT.
Imputing counterfactual outcomes for the treated from data on the control is used in the literature on matrix completion, e.g., Athey et al. (2021); Bai and Ng (2021); Fernández-Val et al. (2021). Our framework has a thin matrix of outcomes. However since we do not observe cross-sectional control units, we cannot appeal to these methods to impute the counterfactuals.
C.2 Simulation: Heterogeneous coefficients
We compare the finite-sample behavior of the Unobserved-Components FAT when the counterfactual process satisfies Assumption 4 with heterogeneous coefficients $\rho_{i}$ and $\delta_{i}$. That is, we consider the same specification of $y_{it}(0)$ as in Section 4.2 with $I_{1}=I_{2}=I_{3}=1$, with the only changes being that $\rho_{i}$ and $\delta_{i}$ vary across individuals.